# Efficient Spectral Element Method for the Euler Equations on Unbounded Domains in Multiple Dimensions

### v1

@@ Abstract
Mitigating the impact of waves leaving a numerical domain has been a persistent challenge in numerical modeling. Reducing wave reflection at the domain boundary is crucial for accurate simulations. Absorbing layers, while common, often incur significant computational costs. This paper introduces an efficient application of a Legendre-Laguerre basis for absorbing layers for two-dimensional non-linear compressible Euler equations. The method couples a spectral-element bounded domain with a semi-infinite region, employing a tensor product of Lagrange and scaled Laguerre basis functions. The semi-infinite region serves as an absorbing layer for our simulations. In comparison to existing methods with similar absorbing layer extensions, our approach, a pioneering application to the Euler equations, demonstrates substantial computational savings. The study marks the first application of semi-infinite elements to mitigate wave reflection in the solution of the Euler equations, particularly in nonhydrostatic atmospheric modeling. A comprehensive set of tests demonstrates the method’s versatility for general systems of conservation laws, with a focus on its effectiveness in damping vertically propagating gravity waves in a linear hydrostatic mountain simulation—a benchmark for atmospheric models. Across all tests, our model consistently exhibits notable performance improvements compared to a traditional Rayleigh damping approach.
@@ 1 Introduction
The second and more common approach is Rayleigh damping, which relaxes one or more of the state variables near the boundary. This approach augments the computational domain by adding additional levels/cells in the form of an absorbing layer that damps outgoing waves or disturbances toward a specified reference (see, e.g. [9, 26, 24, 28].) The effectiveness of these methods often depends on a tunable relaxation coefficient and thickness of the absorbing layer and, while many different profiles have been tested for defining the relaxation coefficient, it is often the case that the absorbing layer’s thickness and function parameters are determined by means of empirical criteria [31]. This approach comes with the significant drawback of having to extend the size of the computational domain outside of the region of interest. The absorbing region must be discretized with a sufficient number of cells or elements to appropriately damp outgoing waves leaving the computational domain. These additional cells or elements will thus incur both an additional computational and communication cost for large, high-resolution, multidimensional problems.
@@
As an alternative to the methods described above, Benacchio and Bonaventura [2] proposed an approach based on scaled Laguerre functions to prevent the reflection of one-dimensional shallow water waves resolved by a spectral collocation approach. This method was subsequently extended by the same authors [3] to work with the discontinuous Galerkin method. In this extension, a standard approximation of the governing equations using Lagrange interpolation is employed in the finite region of the domain, while a scaled Laguerre basis is used in the semi-infinite region. The approach was further applied to solve the advection-diffusion equation on unbounded domains [43], and extended to address hyperbolic and parabolic problems in two dimensions [44]. For the continuous spectral element solution of the Stokes and Navier-Stokes equations governing incompressible flows, a coupled Legendre-Laguerre solution was proposed in [47], building on the earlier work presented in [36, 37].
@@
The current study extends the cited work to the solution of the non-linear Euler equations for compressible flows with gravity effects. Our approach couples a bounded domain discretized using nodal continuous Galerkin spectral elements with tensor product bases, to a semi-infinite domain discretized using nodal continuous Galerkin elements but using a basis that is the tensor product of a Lagrange polynomial basis on Legendre-Gauss-Lobatto (LGL) points in the normal direction to the outgoing flow, and a Laguerre function basis on Laguerre-Gauss-Radau (LGR) points along the direction of the outgoing wave. Like [47, 3, 43, 44], this approach neither seeks to find an analytical expression for the transmissive boundary conditions as done in e.g.[8, 23], nor does it attempt to be a selective damping approach that targets specific wave lengths for damping e.g.[1, 32, 4, 17]. The use of the Laguerre function basis on semi-infinite elements allows an effective approximation of asymptotic decay and lends itself to the construction of efficient and effective absorbing layers that directly approximate the equations at hand; moreover, the use of nodal continuous Galerkin elements with tensor product bases allows us to use computationally efficient algorithms while making the coupling between our finite and semi-infinite elements seamless.
@@
To the authors knowledge, this paper presents the first application of Laguerre-based absorbing layers to the two-dimensional solution of the Euler equations. Due to the importance of effective absorbing layers in atmospheric simulations, in this paper we study their effectiveness to damp outgoing gravity waves although their applicability is not limited to this problem. The study shows that similarly to [44], the Laguerre absorbing layer is many times cheaper than a classical Rayleigh damping approach and that the larger the size of the desired absorbing layer the lower this relative cost tends to be.
@@
The remainder of this paper is organized as follows: §2 presents the governing equations. §3 presents the nodal continuous Galerkin spectral and semi-infinite element method, the algorithms used to construct the solution on semi-infinite elements, and other numerical techniques that are implemented in Jexpresso [30], a new open-source spectral element code written in performant Julia. Numerical results for a hierarchy of 1D and 2D tests are presented in §4 (All of the test cases presented in this paper are documented and ready to run by the user of the code), followed by discussion in §5and conclusions in §6. The pseudo-code to compute the element right-hand side on semi-infinite elements is given in the Appendix.
@@ 2 Governing equations
This paper considers a general system of PDEs, written in conservation form, on the semi-infinite domain $\Omega~{}\forall~{}t=0,...T_{\rm end}$:
$\frac{\partial\textbf{q}}{\partial t}+\frac{\partial\textbf{F}(\textbf{q})}{\partial x}+\frac{\partial\textbf{G}(\textbf{q})}{\partial z}=\textbf{S}(\textbf{q})+\textbf{V}(\textbf{q}).$ (1)
where the state vector q, flux vectors F and G, source vector S, and diffusion vector V are problem-dependent. Those that are not explicitly stated are taken to be zero vectors of the appropriate size:
@@
    ${\bf q}=\begin{bmatrix}u\\ v\end{bmatrix}\quad{\bf F}=\begin{bmatrix}v\\ u\end{bmatrix},$ (2)
@@
    ${\bf q}=\begin{bmatrix}h\\ u\end{bmatrix}\quad{\bf F}=\begin{bmatrix}Uh+Hu\\ gh+Uu\end{bmatrix},$ (3)
@@
    2D Helmholtz:
    ${\bf S}=\begin{bmatrix}\alpha^{2}u+f(x,z)\end{bmatrix}\quad{\bf V}=\begin{bmatrix}u_{xx}+u_{zz}\end{bmatrix},$ (4)
@@
    ${\bf q}=\begin{bmatrix}q\\ \end{bmatrix}\quad{\bf F}=\begin{bmatrix}qu\\ \end{bmatrix}\quad{\bf G}=\begin{bmatrix}qv\\ \end{bmatrix}\quad{\bf V}=\nu\begin{bmatrix}q_{xx}+q_{zz}\end{bmatrix}$ (5)
@@
    ${\bf q}=\begin{bmatrix}\rho\\ \rho u\\ \rho v\\ \rho\theta\end{bmatrix}\quad{\bf F}=\begin{bmatrix}\rho u\\ \rho u^{2}+p\\ \rho uv\\ \rho u\theta\end{bmatrix}\quad{\bf G}=\begin{bmatrix}\rho v\\ \rho vu\\ \rho v^{2}+p\\ \rho v\theta\end{bmatrix}\quad{\bf S}=\begin{bmatrix}0\\ 0\\ -\rho g\\ 0\end{bmatrix}\quad{\bf V}=\mu\begin{bmatrix}0\\ u_{xx}+u_{zz}\\ v_{xx}+v_{zz}\\ \theta_{xx}+\theta_{zz}\end{bmatrix},$ (6)
    where $\rho$ is density, $u$ is horizontal velocity, $v$ is vertical velocity, $\theta$ is potential temperature, $p$ is pressure, and $\mu$ is a constant coefficient of dynamic viscosity used for stabilization purposes.
@@
$\frac{\partial p^{\prime}}{\partial z}=-\rho^{\prime}g,$
where the primed quantities $\alpha^{\prime}=\alpha-\alpha_{0}$ are the perturbations with respect to a hydrostatically balanced background state $\alpha_{0}$ (see e.g. [29] and citations therein). In addition, the proposed semi-infinite element approach requires that each component of the state vector decays to zero as the vertical coordinate $z\rightarrow\infty$. While density decays to zero as $z\rightarrow 0$, neither velocity nor potential temperature approach zero at high altitudes; hence, perturbation variables allows the proposed method to respect this decay condition while maintaining a physically meaningful reference state.
@@ 3 Numerical methods
The governing equations are discretized in space by means of high-order spectral elements with tensor-product Lagrange bases. Given a computational domain $\Omega$, we first subdivide it into a finite domain $\Omega^{F}$ and a semi-infinite domain $\Omega^{S}$. Each of these subdomains are then divided into conforming quadrilaterals that are either finite spectral elements $\Omega^{F}_{e}$, or semi-infinite elements $\Omega^{S}_{e}$ to form the discrete domain $\Omega_{h}$ as
@@
In time, the governing equations are discretized by means of an explicit 3-stage, $3^{rd}$-order Runge-Kutta approximation. Jexpresso [30], the open-source spectral element code utilized for this study, relies on the DifferentialEquation.jl package by [34] for time integration.
@@ 3.1 Basis functions on the finite domain
The Legendre-Gauss-Lobatto (LGL) points are used both as interpolation and integration nodes (inexact integration). The LGL points are zeros the $N$-th degree Lobatto polynomial, given by
@@
where $\boldsymbol{\xi}=(\xi,\eta)$, $\hat{f_{l}}$ are the expansion coefficients, and $\psi_{l}$ are two-dimensional nodal basis functions defined as the tensor product of the 1D Lagrange polynomials
@@ 3.2 Basis functions on the semi-infinite domain
The semi infinite domain uses both a different set of basis functions and integration points than its finite domain counterpart. First, define the Laguerre polynomials using their three term recurrence relation [37, Eq. (3.3)]:
@@
$L_{k}(\xi)=\frac{2k-1-\xi}{k}L_{k-1}(\xi)-\frac{k-1}{k}L_{k-2}(x),\forall k\geq 2.$ (17c)
Additionally the first derivative of the $k$-th Laguerre polynomial is [37, Eq. (3.5)]:
@@
The Laguerre polynomials are orthogonal on the semi-infinite interval $[0,\infty)$ with respect to an exponentially decaying weight. This orthogonality property can be written as follows:
@@
The Laguerre-Gauss-Radau (LGR) points are the roots of $\xi L^{\prime}_{N+1}(\xi)$ for a fixed integer N. The LGR points $\{\xi_{j}\}_{j=0,N}$ will be used to construct a nodal spectral element on the semi-infinite elements. We compute the LGR points using the Eigenvalue Method [15, 37], which forms a tridiagonal matrix using the coefficients in the three-term recurrence relationship (3.2) and solves an eigenvalue problem. This method is stable and robust at very high-order and we have tested it for orders up to 60. A scaling factor $f$ is used to adjust the positions of the LGR nodes, such that the physical nodes on the semi-infinite element become are $\{x\}^{N+1}_{i}=1=f\{\xi\}^{N+1}_{i=1}$. This scaling factor allows us to adjust the size of the semi-infinite element and is then taken into account through the metric term $\partial\eta/\partial z$. Next we construct the Lagrange-Laguerre interpolating polynomials associated with the LGR points $\{\xi_{j}\}_{j=0,\dots,N}$ points following [36]:
@@
We now introduce the scaled Laguerre function (SLF) [36]
$\hat{L}_{i}(\xi)=e^{-\frac{\xi}{2}}L_{i}(\xi).$ (22)
The substitution of (22) into (19) yields:
@@
indicating that the SLFs form an orthogonal basis on $L^{2}(\mathbb{R}^{+})$. Each SLF decays exponentially as $\xi\rightarrow\infty$. This property is illustrated in figure 2, which shows the first six Laguerre functions $\hat{L}_{i}(\xi)$. Thanks to this property, the SLFs (22) are ideal for approximating functions in an absorbing layer; the layer damps any outgoing perturbations by enforcing the exponential decay property.
Remark: We limit the use of damping terms to $\Omega^{S}$ and essentially overlap our sponge layer with the semi-infinite elements. In the results, we will describe the damping coefficients used for each test and we will show that minimal reflection can be obtained while relying on this approach.
Figure 2: First six scaled Laguerre functions (SLFs) specified by (22).

### v3

@@ Abstract
Mitigating the impact of waves leaving a numerical domain has been a persistent challenge in numerical modeling. Reducing wave reflection at the domain boundary is crucial for accurate simulations. Absorbing layers, while common, often incur significant computational costs. This paper introduces an efficient application of a Legendre-Laguerre basis for absorbing layers for two-dimensional non-linear compressible Euler equations. The method couples a spectral-element bounded domain with a semi-infinite region, employing a tensor product of Lagrange and scaled Laguerre basis functions. Semi-infinite elements are used in the absorbing layer with Rayleigh damping. In comparison to existing methods with similar absorbing layer extensions, this approach, a pioneering application to the Euler equations of compressible and stratified flows, demonstrates substantial computational savings. The study marks the first application of semi-infinite elements to mitigate wave reflection in the solution of the Euler equations, particularly in nonhydrostatic atmospheric modeling. A comprehensive set of tests demonstrates the method’s versatility for general systems of conservation laws, with a focus on its effectiveness in damping vertically propagating mountain gravity waves, a benchmark for atmospheric models. Across all tests, the model presented in this paper consistently exhibits notable performance improvements compared to a traditional Rayleigh damping approach.
@@ 1 Introduction
A second and more robust approach is Rayleigh damping, which relaxes one or more of the state variables in an absorbing layer near the boundary. This approach augments the computational domain by adding additional levels/cells in the absorbing layer that damps outgoing waves or disturbances toward a specified reference (see, e.g. [11, 31, 29, 34].) The effectiveness of these methods often depends on a tunable relaxation coefficient and thickness of the absorbing layer. While many profiles have been employed, often the absorbing layer’s thickness and parameters are determined empirically [37]. Since the absorbing region must be discretized with additional cells or elements to appropriately damp outgoing waves, these additional cells/elements incur additional computational and communication cost for large, high-resolution, multidimensional problems.
@@
As a cost-effective alternative to stadard Rayleigh damping, Benacchio and Bonaventura [2] proposed a spectral collocation approach based on scaled Laguerre functions to prevent the reflection of one-dimensional shallow water waves. This method was subsequently extended by the same authors [3] to the discontinuous Galerkin (DG) method. The governing equations are discretized with Lagrange polynomials in the finite region of the domain, while a scaled Laguerre basis is used in the semi-infinite region. The approach was further applied to solve the advection-diffusion equation on unbounded domains [53], and extended to address hyperbolic and parabolic problems in two dimensions [52]. For the continuous spectral element solution of the Stokes and Navier-Stokes equations governing incompressible flows, a coupled Legendre-Laguerre solution was proposed in [56], building on the earlier work presented in [43, 44].
@@
The current study uses the nodal continuous Galerkin (CG) combined with semi-infinite elements to solve the compressible nonlinear Euler equations for with gravity. This paper couples a bounded domain discretized using nodal continuous Galerkin spectral elements with tensor product bases, to a semi-infinite domain discretized using a basis that is the tensor product of a Lagrange polynomial basis on Legendre-Gauss-Lobatto (LGL) points in the normal direction to the outgoing flow, and a Laguerre function basis on Laguerre-Gauss-Radau (LGR) points along the direction of the outgoing wave. Like [56, 3, 53, 52], this approach neither seeks to find an analytical expression for the transmissive boundary conditions as done in e.g.[9, 28], nor does it attempt to be a selective damping approach that targets specific wave lengths for damping e.g.[1, 38, 4, 22]. The use of the Laguerre function basis on semi-infinite elements allows an effective approximation of exponential decay and lends itself to the construction of efficient and effective Rayleigh damping layers. Moreover, the use of nodal continuous Galerkin elements with tensor product bases allows us to use computationally efficient algorithms while making the coupling between finite and semi-infinite elements seamless.
@@
To the authors’ knowledge, this paper presents the first application of Laguerre-based absorbing layers to the two-dimensional solution of the Euler equations of compressible flows. Due to the importance of effective absorbing layers in atmospheric simulations, this paper studies their effectiveness to damp outgoing gravity waves although their applicability is not limited to this problem. As noted by [52], the Laguerre absorbing layer is many times cheaper than a classical Rayleigh damping approach and scales favorably with respect to the absorbing layer thickness.
@@
The remainder of this paper is organized as follows: §2 presents the governing equations. §3 presents the nodal continuous Galerkin spectral and semi-infinite element method, the algorithms used to construct the solution on semi-infinite elements, and other numerical techniques that are implemented in Jexpresso [36], a new open-source spectral element code written in performant Julia. Numerical results for a hierarchy of 1D and 2D tests are presented in §4 (All of the test cases presented in this paper are documented and ready to run by the user of the code), followed by discussion in §5 and conclusions in §6. The pseudo-code to compute the element right-hand side on semi-infinite elements is given in the Appendix A while Appendix B provides a description of how to extend the current method to three dimensions. Appendix C shows that for mountain wave problems, no significant errors are introduced by the method when compared to a standard spectral element approach.
@@ 2 Governing equations
This paper considers a general system of PDEs, written in conservation form, on the semi-infinite domain $\Omega$ for $0\leq t\leq T_{\mbox{end}}$:
$\frac{\partial\textbf{q}}{\partial t}+\frac{\partial\textbf{F}(\textbf{q})}{\partial x}+\frac{\partial\textbf{G}(\textbf{q})}{\partial z}=\textbf{S}(\textbf{q})+\textbf{V}(\textbf{q})+\textbf{R}(\textbf{q}),$ (1)
where the state vector q, flux vectors F and G, source vector S, and diffusion vector V are problem-dependent. A Rayleigh damping layer R is used to damp outgoing waves, where required, is defined as
$\mathbf{R}(\mathbf{q})=-\gamma(z)\left(\mathbf{q}-\mathbf{q}_{0}\right),$ (2)
where $\gamma(z)$ is a damping coefficient to be defined when needed and $\mathbf{q}_{0}$ is the reference state. The following equations are be solved in this paper:
@@
    ${\bf q}=\begin{bmatrix}u\\ v\end{bmatrix},\quad{\bf F}=\begin{bmatrix}v\\ u\end{bmatrix},$ (3)
@@
    ${\bf q}=\begin{bmatrix}h\\ u\end{bmatrix},\quad{\bf F}=\begin{bmatrix}Uh+Hu\\ gh+Uu\end{bmatrix},$ (4)
@@
    2D Helmholtz equation:
    ${\bf S}=\begin{bmatrix}\alpha^{2}u+f(x,z)\end{bmatrix},\quad{\bf V}=\begin{bmatrix}u_{xx}+u_{zz}\end{bmatrix},$ (5)
@@
    ${\bf q}=\begin{bmatrix}q\\ \end{bmatrix},\quad{\bf F}=\begin{bmatrix}qu\\ \end{bmatrix},\quad{\bf G}=\begin{bmatrix}qv\\ \end{bmatrix},\quad{\bf V}=\nu\begin{bmatrix}q_{xx}+q_{zz}\end{bmatrix}.$ (6)
@@
    ${\bf q}=\begin{bmatrix}\rho\\ \rho u\\ \rho v\\ \rho\theta\end{bmatrix},\quad{\bf F}=\begin{bmatrix}\rho u\\ \rho u^{2}+p\\ \rho uv\\ \rho u\theta\end{bmatrix},\quad{\bf G}=\begin{bmatrix}\rho v\\ \rho vu\\ \rho v^{2}+p\\ \rho v\theta\end{bmatrix},\quad{\bf S}=\begin{bmatrix}0\\ 0\\ -\rho g\\ 0\end{bmatrix},\quad{\bf V}=\begin{bmatrix}0\\ \mu(u_{xx}+u_{zz})\\ \mu(v_{xx}+v_{zz})\\ \kappa(\theta_{xx}+\theta_{zz})\end{bmatrix},$ (7)
    where $\rho$ is density, $u$ is horizontal velocity, $v$ is vertical velocity, $\theta$ is potential temperature, $p$ is pressure, and $\mu$ and $\kappa$ are constant artificial viscosity coefficients.
@@
$\frac{\partial p^{\prime}}{\partial z}\quad\rm{and\quad}\rho^{\prime}g,$
where the primed quantities $\alpha^{\prime}=\alpha-\alpha_{0}$ are the perturbations with respect to a hydrostatically balanced background state $\alpha_{0}$ (see e.g. [35] and citations therein). In addition, the proposed semi-infinite element approach requires that each component of the state vector decays to zero as the vertical coordinate $z\rightarrow\infty$. While density decays to zero as $z\rightarrow\infty$, neither velocity nor potential temperature approach zero at high altitudes; hence, perturbation variables allows the proposed method to respect this decay condition while maintaining a physically meaningful reference state.
@@ 3 Numerical methods
The governing equations are discretized in space by means of high-order spectral elements with tensor-product Lagrange bases. In contrast to the modal discontinuous Galerkin approach of [3], we use nodal continuous Galerkin spectral elements. Given a computational domain $\Omega$, we first subdivide it into a finite domain $\Omega^{F}$ and a semi-infinite domain $\Omega^{S}$. Each of these subdomains are then divided into conforming quadrilaterals that are either finite spectral elements $\Omega^{F}_{e}$, or semi-infinite elements $\Omega^{S}_{e}$ to form the discrete domain $\Omega_{h}$ as
@@
In time, the governing equations are discretized by means of an explicit 3-stage, $3^{rd}$-order Runge-Kutta approximation. Jexpresso [36], the open-source spectral element code utilized for this study, relies on the DifferentialEquation.jl package [40] for time integration.
@@ 3.1 Basis functions on the finite domain
The Legendre-Gauss-Lobatto (LGL) points are used both as interpolation and integration nodes (inexact integration). The LGL points are the zeros of the $N$-th degree Lobatto polynomial, given by
@@
where $\boldsymbol{\xi}=(\xi,\eta)$, $\hat{f_{l}}(t)$ are the expansion coefficients, and $\psi_{l}(\boldsymbol{\xi})$ are two-dimensional nodal basis functions defined as the tensor product of the 1D Lagrange polynomials
@@ 3.2 Basis functions on the semi-infinite domain
The semi infinite domain uses both a different set of basis functions and integration points than its finite domain counterpart. First, define the Laguerre polynomials using their three term recurrence relation [44, Eq. (3.3)]:
@@
$L_{k}(\xi)=\frac{2k-1-\xi}{k}L_{k-1}(\xi)-\frac{k-1}{k}L_{k-2}(\xi),\forall k\geq 2.$ (18c)
Additionally the first derivative of the $k$-th Laguerre polynomial is [44, Eq. (3.5)]:
@@
The Laguerre polynomials are orthogonal on the semi-infinite interval $[0,\infty)$ with respect to an exponentially decaying weight as follows:
@@
The Laguerre-Gauss-Radau (LGR) points are the roots of $\xi L^{\prime}_{N+1}(\xi)$ for a fixed integer N. The LGR points $\{\xi_{j}\}_{j=0,N}$ will be used to construct a nodal spectral element on the semi-infinite elements. We compute the LGR points using the Eigenvalue Method [20, 44], which forms a tridiagonal matrix using the coefficients in the three-term recurrence relationship (18) and solves an eigenvalue problem. This method is stable and robust at very high-order and we have tested it for orders up to 60. Following [43], the scaling factor $\lambda$ adjusts the LGR nodes, such that the physical nodes on the semi-infinite element are $\{x\}^{N+1}_{i=1}=\lambda\{\xi\}^{N+1}_{i=1}$. This scaling factor allows us to adjust the effective length of the semi-infinite element for a given problem.
Next we construct the Lagrange-Laguerre interpolating polynomials associated with the LGR points $\{\xi_{j}\}_{j=0,\dots,N}$ points following [43]:
@@
We now introduce the scaled Laguerre function (SLF) [43, 44, 2]
$\hat{L}_{i}(\xi)=e^{-\frac{\xi}{2\lambda}}L_{i}\left(\frac{\xi}{\lambda}\right),$ (23)
where $\lambda$ is a scaling factor and represents a characteristic length. Note that this notation is equivalent to the notation in [2] for $\lambda=\beta^{-1}$. Applying (20) yields
@@
indicating that the SLFs form an orthogonal basis on $L^{2}(\mathbb{R}^{+})$. Each SLF decays exponentially as $\xi\rightarrow\infty$ for any $\lambda>0$. This property is illustrated in figure 2, which shows the first six Laguerre functions $\hat{L}_{i}(\xi)$ for $\lambda=1$. Thanks to this property, the SLFs (23) are ideal for approximating functions in an absorbing layer; the layer damps any outgoing perturbations by enforcing the exponential decay property.
Remark: We limit the use of damping terms to $\Omega^{S}$ and essentially overlap the sponge layer with the semi-infinite elements. In the results, we will describe the damping coefficients used for each test and we will show that minimal reflection can be obtained while relying on this approach.
Figure 2: First six scaled Laguerre functions (SLFs) specified by (23) with scaling factor $\lambda=1$.